rods»view/81

Status Summary for ROD ID 81

<<Prev | Next>>

Test NameStatusNotesModified
Board Status 1 Some Fifo to V2 channels stuck high. Board reboots when tapped at current limit 4A on both channels; also the lower bank of LED's as well as the load-indicator LED nearest U40 shut off after conn test without obvious provocation; resumed working after reboot -Bill. While in the crate both top and bottom slink up LEDs on the front panel started blinking. -Alex 10-11-2009
loop data 1
Passes with new mig. fails w/ Buffer: 07ffffff Correct: 00201008 for all words.
23-06-2008
slink 1
23-06-2008
dram 1
23-06-2008
nse communication 1
23-06-2008
sram
23-06-2008
optical loop 1
Passes after exchanging the Slink cards. Top and Bottom Fail with 1016 errors each time With New Firmware and delays set to 31, both sides passed
23-06-2008
patch panel 1
Replaced QPLL Crystal and everything works. -AC 6/23/08 Optical Links will not lock reliably and Rocket IO values are Always zero. Check next to TTC Board went Crazy during testing. Unable to connect through the VME. LEDs flashing. Removed from crate.
23-06-2008
qpll lock 1
23-06-2008
ttc connectivity 1
23-06-2008
test fifo 0
V2-testfifo fault.
23-06-2008
ddr reset
23-06-2008
assembly 1
23-06-2008
fuses 1
23-06-2008
U31 1
23-06-2008
U32 1
23-06-2008
U33 1
23-06-2008
U34 1
23-06-2008
U35 1
23-06-2008
U81 1
23-06-2008
U82 1
23-06-2008
I5V_initial 1
23-06-2008
I3V initial 1
23-06-2008
proms load 1
23-06-2008
I5V proms 1
23-06-2008
I3V proms 1
23-06-2008
V4T temp 1
23-06-2008
V4B temp 1
23-06-2008
connectivity 1
23-06-2008